UT 20150430 - 20150503 LUCI2 Closed-dome ACF & repro-tests

Observer: APramskiy, JHeidt, DThompson (remote from Tucson)
Telescope Operator: various
LBTO Support: TSargent

Details

April 30 2015

APramskiy & JHeidt en route to Tucson, DThompson wanted to perform some repro-tests

However, he faced an issue with the Camera-wheel (IT #5553), ie he couldn't set the
camera position properly. To solve this an intervention might be requred but is currently under investigation

May 1 2015

No tests, some discussions on the outline of the upcoming run.

JHeidt brought a frosted glass to Steward which was delivered to the LBT the same day

May 2 2015

ACF-tests using the srre-mode (sample-up-the-ramp expanded). Not successfull, holes required
for ACFcould not be seen on the subimages. Various tests were performed to find the reason, but failed

The frosted screen has to be added to the exit of the illuminating sphere inside the Calibration Unit
of LUCI2 by JMorris

May 3 2015

Files are luci2.20150503.NNNN.fits

Dthompson discussed with RMathar the issue with the srre-mode. It is not suitable
since it does an reset before each sample with the difference is being saved).
The recommendation is to use the standard srr-mode. First tests indicate that this is appropriate

### Check on whether the Calibration Unit no homogeneoulsy illuminates

Initialize all units...
Calibration unit takes a long time. It was IN the beam when I initialized. Init moved it OUT. LUCI2 GUI shows the CU as "out"
but the "IN" checkbox is selected. Refresh does not change this. Two problems: INIT should leave it wherever it is, just confirm
that it is initialized. Also, the "IN" radio button on the LUCI2 GUI remained selected even after a refresh and the status window
reporting "OUT".

0087 1 x 2.75s Ks lamp off image
0088 1 x 2.75s Ks Halo3 lamp on image

0089 1 x 2.75s H2 lamp off image
0090 1 x 2.75s H2 Halo2 lamp on image

=> The gradient across the detector seen is now reduced from about 60% to 20-25%,
BUT the gradient remains. Moreover, the gradient is now point symmetric with respect
to the one seen before the frosted screen has been inserted. In addition, there is some
vignetting at the top of the images.

flatH2_lamp2.png flatH2_lamp2_withdiffuser.png

Left) Flat H2/Halo2 without frosted glass. Right) Flat H2/Halo2 with frosted glass. The dynamic range with which the images
are shown is in both cases the same.

### Filter Wheel reproducibility
FW1_repro_270.xml - error...rotator angle not sent to TCS
Tom said "commit" may not have been sent

Running generic script manually...
Set ROT = 270 on MCS GUI
FW1_repro.xml 0091 0094
Error during commit - java null pointer exception
commit again on OEP...still looks like it hangs
reset, refresh, find head...@ beginning of script now
commit...did nothing...commit again...put in filter and stopped, still problems
Stop scheduler from LMC, stop OEP, start scheduler, start OEP
reset, refresh, find head...@ beginning of script now

FW1_repro.xml 0095 0104
Error committing compensation mirror setup, but script continues (before 0105)...
Script ends with moving CU out, but it should not (this is inserted by scriptor)

Queue Actions...the whole import of a program needs to be modified...
Should be able to clear visible observing prgram list to speed things up
Imported program list should be displayed from newest @ top to oldest @ bottom
Newly imported program then appears @ top
Clarify information presented (coordinate w/scriptor updates as needed)

OEP...information presented in the individual items needs to be clarified and made more compact if possible

FW2_repro.xml 0105 0113
Edited script to remove last item
Commit...finished without problems
One or more images saturated?

Set ROT = 180
FW2_repro.xml 0114 0123
Commit...all items executed
One or more images saturated?

FW1_repro.xml 0124 0133
Modified script to remove the last item that moves the CU out and turned off lamps
Commit...all items executed
Last item shows "running" when all items processed...AP says this is a known bug

Tom says the NATIVE_ANGLE bug is fixed...

FW1_repro_090.xml 0134 0143
ROT angle went to 90.0
Commit...all items executed

FW2_repro_090.xml 0144 0153
Commit...all items executed

FW1_repro_000.xml 0154
Just pause after 1st exposure...test for TSargent

The FW2_repro* scripts use halo2 and broad filters, most data saturated

DT Test for Sensitivity

0155 @ default position for J_high filter ~56000ADU in 3.0 sec w/Halo2

0156 Lamp off image with Pbeta (1.282um)
0157 Halo2 ON with Pbeta at nominal position ~8000 ADU
On MCU GUI
FW1 tab
Turn on holding current
Set counter to 0
Move FW1 +1 step
0158 +1 step
0159 +2 steps
0160 +3
0161 +4
0162 +5 steps
On MCU GUI
Holding current off
On LUCI2 GUI
Initialize FW1
0163 back at nominal position

Move 'N30 LS neu 1' to FPU for AFC testing
MOS error moving in transport to the open position
Initialize MOS unit successful
Mask to FPU successful

-- %USERSIG{JochenHeidt - 2015-05-03}%
I Attachment Action Size Date Who Comment
flatH2_lamp2.pngpng flatH2_lamp2.png manage 498 K 03 May 2015 - 22:23 JochenHeidt Flat H2/Halo 2 without frosted glass
flatH2_lamp2_withdiffuser.pngpng flatH2_lamp2_withdiffuser.png manage 527 K 03 May 2015 - 22:24 JochenHeidt Flat H2/Halo 2 with frosted glass
Topic revision: r1 - 03 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback